A pharmacist receives a shipment of 25 bottles of a drug and has 4 of the bottles tested. If 5 of the 25 bottles are contaminated, what is the probability that at least 3 of the tested bottles are contaminated? Express your answer as a fraction or a decimal number rounded to four decimal places.
Answer #1

P[ bottle is contaminated ] = 5/25

P[ bottle is contaminated ] = 0.2

Sample size = n = 4

The population proportion of success is p = 0.2, also, 1 - p = 1 - 0.2 = 0.8, and the sample size is n= 4. We need to compute

This implies that